OpenStudy (anonymous):

Write the fractions in order from smallest to largest. 7/10,3/20,22/25,2/5

OpenStudy (anonymous):

2/5,22/25,7/10,3/20

Parth (parthkohli):

No rohan, it's wrong!

OpenStudy (anonymous):

how do you figure that out? and thank you

OpenStudy (anonymous):

never mind i was sleeping!

Parth (parthkohli):

3/20, 2/5, 7/10, 22/25 Get the denominators same and then see which one has the highest numerator and denominator. You get the answer.

Parth (parthkohli):

3/20 = 15/100 2/5 = 40/100 7/10 = 70/100 22/25 = 88/100 Now, can you solve it? :D

OpenStudy (anonymous):

oh ok

Parth (parthkohli):

Can I get a medal please?

OpenStudy (anonymous):

ok i entered them and it says its wrong only 3/20 is right?

OpenStudy (anonymous):

Im thinking its my test that is not accepting the answer!

Parth (parthkohli):

First get the denominators same, then check them, them get them back to the simplest form.

OpenStudy (anonymous):

ok gv me a min thx

Parth (parthkohli):

3/20, 2/5, 7/10, 22/25 this is the answer after simplifying back.

OpenStudy (anonymous):

yes thats what i got, I really think its my online app, thank you and this is the first time if been on this website how do i gv you a medal? lol

OpenStudy (anonymous):

I think I just figured it out haha